Breaking even occurs when the revenues for a buisness equal the cost. A local children's museum studied their costs and revvenues from paid admission. They found that their break even time is given by the equation of 2h^2-2h-24=0 where h is the number of hours the museum is open per day. How many hours must the museum be open per day to reach the break even point
Answer by stanbon(75887)

They found that their break even time is given by the equation of 2h^2-2h-24=0 where h is the number of hours the museum is open per day. How many hours must the museum be open per day to reach the break even point
-------
2(h^2-h-12) = 0
--------------------
(h-4)(x+3) = 0
---
Positive solution:
h = 4 hours
--------------------
Cheers,
Stan H.
=================